# fixedprepend, caption=FIXED, color=green

### v2

@@ Abstract
While large language models (LLMs) have achieved impressive progress, their application in scientific domains such as chemistry remains hindered by shallow domain understanding and limited reasoning capabilities. In this work, we focus on the specific field of chemistry and develop a Chemical Reasoner LLM, ChemDFM-R [^1] The inference code and the model parameters will be open-sourced.. We first construct a comprehensive dataset of atomized knowledge points to enhance the model’s understanding of the fundamental principles and logical structure of chemistry. Then, we propose a mix-sourced distillation strategy that integrates expert-curated knowledge with general-domain reasoning skills, followed by domain-specific reinforcement learning to enhance chemical reasoning. Experiments on diverse chemical benchmarks demonstrate that ChemDFM-R achieves cutting-edge performance while providing interpretable, rationale-driven outputs. Further case studies illustrate how explicit reasoning chains significantly improve the reliability, transparency, and practical utility of the model in real-world human-AI collaboration scenarios.
@@ 1 Introduction
With the remarkable capabilities and performance demonstrated by large language models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Team et al., 2023), the development of domain-specialized LLMs has emerged as a popular approach to addressing complex problems (Hendrycks et al., 2020; Chen et al., 2021; Wang et al., 2024). Specifically, in the scientific domain, many efforts focus on specializing LLMs for specific tasks (Yu et al., 2024; Antunes et al., 2024; Sriram et al., 2024) or building general-purpose scientific assistants (Zhao et al., 2025b; Zhang et al., 2024a; Zhao et al., 2024; Zhang et al., 2025b; Tan et al., 2025). However, given the inherent complexity and high demand for reliability in the scientific domain, current models often struggle with inadequate performance and limited interpretability, which significantly hinders their practical applicability.
@@
Recently, great success has been achieved in constructing reasoner LLMs in the general domain (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025; Comanici et al., 2025). Beyond enhancing overall model performance, the reasoning-before-answering pattern directly exhibits how and why the LLM comes up to this answer and therefore markedly improves the reliability and interpretability of the LLM’s answer. Through the generated rationale, people can confirm the correctness of the answer or identify why the model makes mistakes. Therefore, reasoning-augmented LLMs offer a promising new approach to addressing the aforementioned challenges in scientific-domain LLMs, potentially enhancing their practical utility.
@@
However, current research on reasoner LLMs has predominantly focused on general domains, as well as mathematics (Yang et al., 2024b; Shao et al., 2024) and programming (Zhu et al., 2024; Hui et al., 2024). In scientific domains where professional knowledge and specific logic are vital and can hardly be harvested from natural language content, the reasoning capability of existing LLMs remains highly limited. There are two reasons that hinder current LLMs from excelling in science-domain-specific reasoning: 1) The understanding of domain knowledge remains superficial as advanced domain knowledge is typically insufficient in general-purpose corpora. Moreover, it is noticed that even current domain-specific LLMs also suffer from shallow understanding of domain knowledge. The literature and high-level phenomena used in current domain LLM training (Taylor et al., 2022; Xie et al., 2023; Zhao et al., 2025b; Zhang et al., 2024a; Yu et al., 2024) often fail to conduct a proper breakdown into internal mechanisms and atomized knowledge points. For example, in the field of chemistry, the types and positions of functional groups within molecules fundamentally determine molecules’ properties and reactivities. However, instead of training on the functional-group-level knowledge, current chemical LLMs usually directly learn the molecule-level knowledge about properties and reactivities. The lack of atomized knowledge significantly constrains the capacity of these models for providing high-quality rationales. 2) The intrinsic reasoning logic in these domains differs significantly from that in general domains such as mathematics and programming, making it difficult for models to generalize relevant reasoning skills through training in general domains. Therefore, specific reasoning training in the vertical domains is needed.
@@
In this work, we focus on addressing the two aforementioned challenges in the chemistry domain and develop a chemical reasoner LLM, ChemDFM-R. Specifically, we consider the presence of functional groups in molecules and the changes of functional groups during reactions to be atomized knowledge points. We develop a toolkit to identify these features from molecules and reactions and incorporate them into the domain pretraining corpus. The resulting corpus contains over 101 billion tokens from 12 million literature, 30 million molecules, and 7 million reactions. To tackle the second issue, we develop a mix-sourced distillation process that can take advantage of both the expertise in the carefully curated knowledge points and the advanced reasoning capabilities of general LLMs. Domain-specific reinforcement learning is utilized following the distillation process to further enhance the reasoning capability. The resulting model, ChemDFM-R, shows promising reasoning capabilities as well as advanced performance in multiple chemical benchmarks and can provide high-quality rationale, helping researchers deeply understand and verify its answer. In short, the contributions of this work are threefold:
@@
  • We developed a toolkit to identify the functional groups of molecules and the changes of functional groups during reactions. With the help of this toolkit, we constructed a 101 billion-token domain pretraining corpus, named ChemFG, featuring atomized functional-group-level knowledge in chemistry.
  • With the base model domain-pretrained with ChemFG from Qwen2.5-14B, we further conducted instruction tuning, mix-sourced distillation, and domain-specific reinforcement learning in order. The proposed distillation and reinforcement learning significantly boost the model’s capability of chemical reasoning and produce the first chemical reasoner LLM, ChemDFM-R.
@@
  • Extensive experiments are conducted, and the results demonstrate the promising chemical reasoning capability of ChemDFM-R. The created model achieves outstanding performance and manages to generate clear and rational rationales, which significantly boost the reliability and interpretability of the final answer.
@@ 2 ChemFG
In the field of chemistry, functional groups serve as the bridge between molecular structures, properties, and reactivities, making them one of the most critical intermediate reasoning steps in chemical reasoning. However, existing training corpora of LLMs often lack detailed information on molecular functional groups, preventing models from directly and precisely learning this atomized chemical knowledge. Therefore, we crawl and construct a functional-group-centered domain pertaining corpus, ChemFG, which contains over 101 billion tokens constructed from 12 million literature, 30 million molecules, and 7 million reactions.
2.1 Raw Data Collection
ChemFG mainly consists of data from three sources: literature, molecules, and reactions.
Literature.
Literature, including papers and textbooks, contains not only the widely accepted chemical knowledge and laws, but also the cutting-edge research in the field of chemistry. Therefore, to take full advantage of chemical literature, we collected over 12 million literature from the open Internet dated prior to January 2022. After further cleaning and deduplication, 79B tokens are obtained from it.
Molecules.
Molecules are the fundamental participants in various chemical processes. Therefore, it is crucial for Chemical LLMs to understand molecular structures and properties. We manage to acquire large-scale molecule datasets from PubChem [^2] https://pubchem.ncbi.nlm.nih.gov/, one of the biggest open accessible chemical databases with more than 100M compounds. We include 30 million molecules along with their notations, descriptions (if applicable), and properties. Besides PubChem, we also leverage the PubChemQC (Nakata and Shimazaki, 2017) dataset, which contains the quantum chemical calculation results of 86M molecules from PubChem, to supplement the quantum chemical properties of these molecules, such as dipole moment and orbital energy. To diversify the final data entry, we randomize the order of the properties and use three different formats: markdown list, markdown tables, and JSON dictionary to formulate the molecule data.
Reactions.
Reactions are the major process in the chemical world. The reactions and non-reactions between molecules are the primary ways in which chemistry impacts our daily lives. In ChemFG, we use the reactions from USPTO-FULL (Dai et al., 2019), one of the most comprehensive open-sourced chemical reaction databases. To avoid data leakage, we exclude the test set of USPTO-FULL, USPTO-MIT (Jin et al., 2017), and USPTO-50K (Schneider et al., 2016) according to the products of reactions. Moreover, to further enhance the data diversity, we leverage the SMILES augmentation method introduced in R-SMILES (Zhong et al., 2022) and achieve a total of 10 times augmentation of data. Finally, a corpus of 7 million reactions is obtained.
@@ 2.2 Functional Group Identification
Although we have collected web-scale molecules and reactions from the open Internet, there are no existing benchmarks that describe the correspondence between functional groups with molecules or reactions. To tackle this issue, we develop a functional group identification toolkit based on thermo library [^3] https://thermo.readthedocs.io/, SMARTS (SMiles ARbitrary Target Specification) notations [^4] https://www.daylight.com/dayhtml/doc/theory/theory.smarts.html, and atom mapping annotations of reactions.
Specifically, we extend and refine the SMARTS list of functional groups provided by thermo to identify the functional groups present in a molecule. After our improvement, the tool can successfully identify 241 types of different common functional groups (see Appendix A), while the original thermo can only identify 83 types with errors. With this functional group identifier, we annotate the functional groups of all the molecules in our domain-pertaining corpus and incorporate them as one of the molecular properties during the formulation of the final data entry.
As for reactions, we annotate the changes of functional groups during reactions with the following process. First, with the help of atom mapping annotations provided by the USPTO-FULL dataset, we identify the reaction centers as the atoms that are involved in bond changes during reactions. Based on these reaction centers, we identify the reacting functional groups as the reactants’ functional groups that contain any of the reaction centers but do not appear in the product. Similarly, the resulting functional groups are defined as the product’s functional groups that contain any of the reaction centers but do not appear in any of the reactants. Finally, the reaction can be described as a reaction between the reacting functional groups that results in the formation of new functional groups. Besides functional groups, there are other structural changes during reactions that are equally important, including ring breaking, ring forming, and bond changes outside functional groups. Therefore, we further construct tools to identify these changes.
@@ 2.3 Quality Control
To ensure the annotation quality of functional groups, we hire three chemical experts to conduct manual inspections. Firstly, all the experts agree that our SMARTS list has already covered the most common functional groups. For molecules, our tool’s annotation accuracy of 100 random samples reaches over 90%, with errors primarily due to corner cases such as rare functional groups or complex interactions between functional groups and aromatic rings. For the annotation of reactions, our tool achieves over 80% accuracy when tested with 100 random samples. The errors mainly arise from invalid reactions or wrong atom mapping annotations.
@@ 3 ChemDFM-R
As outlined in Figure 1, the training pipeline can be divided into two parts: 1) Domain Pretraining and Instruction Tuning (§˜3.1), where the basic general LLM is trained with atomized chemical knowledge; 2) Distillation and Reinforcement Learning (§˜3.2), where the model’s chemical reasoning capability is enhanced.
@@ 3.1 Atomized Chemical Knowledge Enhancement
In domain pretraining, we leverage the 101 billion-token ChemFG corpus introduced in Section 2 to familiarize our model with the knowledge related to functional groups. Considering that general knowledge is also vital for Chemical LLMs, we train our model from one of the most advanced general LLMs, QWen2.5-14B (Yang et al., 2024a). We also incorporate a substantial amount of general-domain pretraining data into our domain pretraining corpus to ensure that the model retains its general capabilities as much as possible.
@@
We believe that the primary goal of instruction tuning is to teach the model how to analyze the purpose and requirements of a given task and establish connections with the knowledge it learned during the pretraining phase. However, previous instruction tuning datasets in the field of chemistry usually suffer from a severe lack of diversity. Existing datasets typically derive their data from predefined chemistry tasks. On the one hand, the variety of these tasks is inherently limited. On the other hand, within the same task, the instructions used to describe the task are highly repetitive. Therefore, we construct a new instruction tuning dataset for ChemDFM-R. Building upon the instruction tuning dataset of ChemDFM, we significantly enhance the diversity of instructions for each task and introduce numerous new chemistry-related tasks, such as scientific paper QA, chemical property ordering, and reaction step prediction. This expansion improves the overall diversity of the dataset. For detailed information on the construction and composition of the instruction tuning dataset, please refer to Appendix B.
To maintain the general capabilities of our model, we mixed our chemical instruction tuning data with general instruction tuning data in a 1:2 ratio. The Qwen2.5-14B model is finetuned for 2 epochs on this mixed dataset after domain pretraining, resulting in the ChemDFM-I model.
@@ 3.2 Chemical Rationale Learning
Figure 2: Comparison of rationales generated by o3-mini with and without additional input information. We mark the correct analyses in the rationale as green, while the wrong ones as red. For more examples and detailed analyses, please refer to Appendix C.
@@
The common distillation method for LLMs involves gathering rationales from an advanced reasoning LLM, such as DeepSeek-R1 (Guo et al., 2025) and o3-mini (OpenAI, 2025b), and training the student model using the same loss as instruction tuning. However, this process assumes that the teacher LLM is capable of generating sufficiently reasonable rationales, which is often false in the chemical field. Owing to the limited understanding of atomized knowledge and chemical logic, even a powerful general-domain reasoner model will highly probably fail to generate accurate and in-depth enough reasoning for chemical problems.
To tackle this problem, we develop a mix-sourced distillation process to take advantage of both the expertise in the carefully curated knowledge points and the advanced reasoning capabilities of general LLMs. Specifically, the entries in the distillation dataset come from three sources: 1) the instruction tuning dataset of ChemDFM-R ($\sim$70%); 2) constructed pseudo-reasoning describing the functional groups of the involved molecules or reactions ($\sim$22%); 3) teachers’ rationales where additional measure is taken to ensure high quality ($\sim$8%).
Moreover, instead of asking teacher models to generate rationales from scratch, we provide them with rich additional information, including the ground truth answer of the problem and the information of the functional groups of the molecules and reactions that appear in the problem. Comparisons of the rationales generated by DeepSeek-R1 with and without the additional information are illustrated in Figure 2, where the rationales generated with full additional information are significantly more valid and in-depth than the other two. More examples and detailed analysis of the generated rationales using different methods are illustrated in Appendix C, which demonstrates the advantages of our method in terms of generating accurate, complete, and logical rationales.
Specifically, we use DeepSeek-R1 and o3-mini as the teacher models. To further validate the quality of the teachers’ rationales, we hired three chemical experts to manually inspect the rationales. Results show that among the sampled 264 rationales, 177 of them (67%) exhibit sufficiently high quality, 60 of them (23%) have minor, acceptable flaws such as reasoning step skipping or missing possibilities, and 27 of them (10%) contain substantive issues such as logic errors or nonsense reasoning. Considering that reinforcement learning will be used after distillation to correct errors and improve performance, and that it is difficult to systematically distinguish these problems in the rationales, we used all the data during the distillation process.
Similar to instruction tuning, we mix our mix-sourced distillation dataset with general data in a 1:2 ratio. Moreover, the general data is also sampled from multiple sources, where $\sim$92% of the entries are sampled from the general data for instruction tuning of ChemDFM-R and $\sim$8% of the entries are sampled from AM-Deepseek-R1-Distill-1.4M (Zhao et al., 2025a). The ChemDFM-I model is finetuned for one epoch on this mixed dataset.
@@
After distillation, reinforcement learning (RL) is leveraged to further enhance the reasoning capabilities of our model. The composition of the RL dataset is illustrated in Figure 1, while more details of our reinforcement learning process are illustrated in Appendix D.

### v3

@@ Abstract
While large language models (LLMs) have achieved impressive progress, their application in scientific domains such as chemistry remains hindered by shallow domain understanding and limited reasoning capabilities. In this work, we focus on the specific field of chemistry and develop a Chemical Reasoning LLM, ChemDFM-R [^1] https://huggingface.co/OpenDFM/ChemDFM-R-14B. We first construct a comprehensive dataset of atomized chemical knowledge, ChemFG, annotating the presence of functional groups in molecules and the changes of functional groups during chemical reactions, to enhance the model’s understanding of the fundamental principles and internal logic of chemistry. Then, we propose a mixed-source distillation method that integrates expertise in atomized knowledge with general reasoning skills, followed by domain-specific reinforcement learning to enhance chemical reasoning. Experiments on diverse chemical benchmarks demonstrate that ChemDFM-R achieves cutting-edge performance while providing interpretable, rationale-driven outputs. Further case studies illustrate how explicit reasoning chains significantly improve the model’s reliability, transparency, and practicality in real-world human-AI collaboration scenarios.
@@ 1 Introduction
With the remarkable capabilities and performance demonstrated by large language models (LLMs) (brown2020language; achiam2023gpt; team2023gemini), the development of domain-specialized LLMs has emerged as a popular approach to addressing complex problems (hendrycks2020measuring; chen2021evaluating; wang2024mmlu). Specifically, many efforts focus on specializing LLMs into scientific domains to build general-purpose scientific assistants (zhao2025chemdfm; zhang2024chemllm; zhao2024chemdfmx; zhang2025large; tan2025chemmllm) which could assist scientists during their research through conversational human-AI interactions. However, given the inherent complexity and high demand for reliability in the scientific domain, current models often struggle with inadequate performance and limited interpretability, which significantly hinders their practicality.
@@
Recently, great success has been achieved in constructing reasoning LLMs in the general domain (jaech2024openai; guo2025deepseek; team2025kimi; comanici2025gemini). Beyond enhancing overall model performance, the reasoning-before-answering pattern directly demonstrates how and why the LLM arrives at the answer, thereby markedly improving the reliability and interpretability of the LLM’s response. Through the generated rationale, people can confirm the correctness of the answer or identify why the model makes mistakes. Therefore, reasoning-augmented LLMs offer a new promising approach to addressing the aforementioned challenges in scientific-domain LLMs, potentially enhancing their practical utility.
@@
Currently, the research on reasoning LLMs has predominantly focused on general domains, such as mathematics (yang2024qwen25math; shao2024deepseekmath) and programming (zhu2024deepseekcoderv2; hui2024qwen25coder). In contrast, the reasoning capability of existing LLMs remains highly limited in scientific domains. There are two reasons that hinder current LLMs from excelling in scientific-domain reasoning. 1) The understanding of domain knowledge remains superficial owing to the shortage of in-depth enough training data. The advanced domain knowledge is typically insufficient in general-purpose corpora, while even in current domain-specific corpora, the domain knowledge is still shallow. Current domain-specific corpora (taylor2022galactica; xie2023darwin; zhao2025chemdfm; zhang2024chemllm; yu2024llasmol) often focus on literature text and high-level phenomena and fail to conduct a proper breakdown into internal mechanisms and atomized knowledge points. For example, in the field of chemistry, the types and positions of functional groups [^2] A group of atoms in a molecule with distinctive chemical properties. Please refer to https://en.wikipedia.org/wiki/Functional_group for more details. within molecules fundamentally determine molecules’ properties and reactivities. However, instead of training on functional-group-level knowledge, current chemical LLMs usually directly learn molecule-level knowledge about properties and reactivities. The lack of atomized knowledge significantly constrains the capacity of these models for providing high-quality rationales. 2) The intrinsic reasoning logic in these domains differs significantly from that in mathematics and programming, making it difficult for models to generalize relevant reasoning skills through training in general domains. This not only prevents general-domain LLMs from performing high-quality reasoning in scientific domains, but also introduces additional challenges for building domain-specific reasoning models, which usually involve distillation before reinforcement learning. The common distillation method involves gathering rationales from advanced reasoning LLMs, such as DeepSeek-R1 (guo2025deepseek) and o3-mini (openai_o3mini), and training student models using supervised finetuning. This process assumes that the teacher LLM is capable of generating sufficiently reasonable rationales. However, this assumption often fails to hold in the scientific fields, such as chemistry. Owing to the limited understanding of atomized knowledge and chemical logic, even powerful general-domain reasoning models will highly probably fail to generate accurate and in-depth reasoning for chemical problems.
@@
In this work, we focus on addressing the two aforementioned challenges in the chemistry domain and develop a chemical reasoning LLM, ChemDFM-R. Specifically, we consider the presence of functional groups in molecules and the changes of functional groups during reactions to be atomized knowledge points. We develop a toolkit to identify these features from molecules and reactions and incorporate them into the domain pretraining corpus. The resulting corpus contains over 101 billion tokens from 12 million literature, 30 million molecules, and 7 million reactions. To equip the model with the capability of chemistry-specific reasoning, we develop a mixed-source distillation process that can take advantage of both the expertise in the carefully curated knowledge points and the advanced reasoning capabilities of general LLMs. Domain-specific reinforcement learning is applied after the distillation process to integrate different capabilities the model learned from corresponding sources and further enhance the reasoning capability. The resulting model, ChemDFM-R, shows promising reasoning capabilities as well as advanced performance in multiple chemical benchmarks and can provide high-quality rationale, helping researchers deeply understand and verify its answer. In short, the contributions of this work are threefold:
@@
  • We developed a toolkit to identify the functional groups of molecules and the changes of functional groups during reactions. Using this toolkit, we built a 101 billion[-]-token domain pretraining corpus, ChemFG, which encodes atomized functional-group–level chemical knowledge.
  • We proposed a mixed-source distillation method to effectively initialize the model’s reasoning ability under limited resources by incorporating both expertise in functional-group knowledge and distilled rationales. With subsequent reinforcement learning, we achieve a chemical reasoning LLM, ChemDFM-R.
@@
  • Extensive experiments demonstrate the promising chemical reasoning capability of ChemDFM-R. The created model achieves outstanding performance and manages to generate clear and rational rationales, which significantly boost the reliability and interpretability of the final answer.
@@ 2 ChemFG
Figure 1: The influence of functional groups
Table 1: Data composition of ChemFG.
Type Sources #Entries #Tokens
Literature Internet 12M 79B
Molecules PubChem & PubChemQC 30M 22B
Reactions USPTO-FULL 7M 0.6B
In the field of chemistry, functional groups serve as the bridge between molecular structures, properties, and reactivities, making them one of the most critical intermediate reasoning steps in chemical reasoning. As demonstrated in Figure 1, functional groups directly influence the properties of molecules and determine the types of reactions that can take place. However, existing training corpora of LLMs often lack detailed information on molecular functional groups, preventing models from directly and precisely learning this atomized chemical knowledge. Therefore, we collect a functional-group-centered domain pretraining corpus, ChemFG, which consists of data from three sources: literature, molecules, and reactions. The basic statistics of ChemFG are shown in Table 1 with details provided in Appendix A.1.
@@ 2.1 Functional Group Identification
Despite the Internet-scale publicly available molecule and reaction corpora, there are no existing databases that describe the correspondence between functional groups and molecules or reactions. To tackle this issue, we develop a functional group identification toolkit based on thermo library [^3] https://thermo.readthedocs.io/ by extending its embedded SMARTS [^4] https://www.daylight.com/dayhtml/doc/theory/theory.smarts.html, SMiles ARbitrary Target Specification, a normalized form of SMILES. list from 83 types of functional groups to 241 and improving its algorithms. With the help of the developed toolkit, we annotate the functional groups of all our domain-pretraining molecule data. Further details are provided in Appendix A.2.
As for reactions, we annotate the changes of functional groups during reactions with the following process. First, with the help of atom mapping annotations provided by the USPTO-FULL dataset, we identify the reaction centers as the atoms that are involved in bond changes during reactions. Based on these reaction centers and our functional group identification toolkit, we identify the functional groups of the reactants that directly participate in the reaction and those of the product that directly result from the reaction. Finally, the reaction can be described as a functional group transformation, where reacting functional groups are converted into product functional groups. Besides functional groups, there are other structural changes during reactions that are equally important, including ring breaking, ring forming, and bond changes outside functional groups. Therefore, we also construct tools to identify these changes in a similar manner.
@@ 2.2 Quality Control
To ensure the annotation quality of functional groups, we hire three graduate-level chemical experts to conduct manual inspections. Firstly, all the experts agree that the extended SMARTS list has already covered the most common functional groups. For molecules, our tool’s annotation accuracy of 100 random samples reaches 98%, with errors primarily due to corner cases such as rare functional groups or complex interactions between functional groups and aromatic rings. For the annotation of reactions, our tool achieves 89% accuracy when tested with 100 random samples. The errors mainly arise from invalid reactions or wrong atom mapping annotations. Examples and analyses of the error are demonstrated in Appendix A.3.
@@ 3 ChemDFM-R
As outlined in Figure 2, the training pipeline of ChemDFM-R can be divided into two parts: 1) Domain Pretraining and Instruction Tuning (§˜3.1), where the basic general LLM is trained with atomized chemical knowledge; 2) Distillation and Reinforcement Learning (§˜3.2), where the model’s chemical reasoning capability is enhanced.
@@ 3.1 Atomized Chemical Knowledge Enhancement
In domain pretraining, we leverage the 101 billion[-]-token ChemFG corpus introduced in Section 2 to familiarize our model with the knowledge related to functional groups. We train our model from one of the most advanced general LLMs, Qwen2.5-14B (yang2024qwen25). Considering that general knowledge is also vital for Chemical LLMs, we also incorporate a substantial amount of general-domain pretraining data into our domain pretraining corpus to ensure that the model retains its general capabilities as much as possible.
@@
The primary goal of instruction tuning is to teach the model how to analyze the purpose and requirements of a given task and make proper use of the knowledge learned in the pretraining phase. However, existing instruction tuning datasets in the field of chemistry are typically derived from well-studied chemistry tasks and suffer from a severe lack of diversity of both task varieties and instruction expressions. Therefore, we construct a new instruction tuning dataset for ChemDFM-R based on the instruction tuning dataset of ChemDFM (zhao2025chemdfm). To improve the overall task and instruction diversity, we introduce numerous new chemistry-related tasks, such as scientific paper QA, chemical property ordering, and reaction step prediction, and perform instruction-rewriting to achieve an average instruction-entry ratio of 1:50. For detailed information on the construction and composition of the instruction tuning dataset, please refer to Appendix B. To maintain the general capabilities of our model, we mixed our chemical instruction tuning data with general instruction tuning data in a 1:2 ratio. The Qwen2.5-14B model is finetuned for 2 epochs on this mixed dataset after domain pretraining, resulting in the ChemDFM-I model.
@@ 3.2 Chemical Rationale Learning
Figure 3: Comparison of rationales generated by o3-mini with and without additional input information. We mark the correct analyses in the rationale as green, while the wrong ones as red. For more examples and detailed analyses, please refer to Appendix C.1.
@@
Specifically, the entries in the distillation dataset come from three sources, each of which corresponds to part of the abilities required for chemical reasoning: 1) the instruction tuning dataset of ChemDFM-R ($\sim$70%) to maintain basic chemical knowledge and prevent catastrophic forgetting; 2) pseudo-reasoning data describing the functional groups of involved molecules or reactions ($\sim$22%) and highlighting vital intermediate reasoning steps and functional group analyses; 3) teachers’ rationales from DeepSeek-R1 and o3-mini ($\sim$8%) which introduce general reasoning patterns to the model and initiate its reasoning capabilities.
To improve the quality and efficiency of the teacher’s rationales in the chemical domain, instead of asking teacher models to generate rationales from scratch, we provide them with rich additional information. Specifically, the teacher models are provided with the question instruction, the ground truth answer, and the functional group information of the molecules and reactions in the question. Comparison of the rationales generated by DeepSeek-R1 with and without the additional information is illustrated in Figure 3. The rationales generated with full additional information are significantly more valid and in-depth than the other two. More examples with detailed analysis of the rationale generation are given in Appendix C.1. Moreover, to comprehensively assess the quality of our distilled rationales, we sampled a small subset of them and conducted a quantitative human assessment. The results are illustrated in Appendix C.2.
Similar to instruction tuning, we mix our mixed-source distillation dataset with general data in a 1:2 ratio. The general data are also sampled from multiple sources, where $\sim$92% of the entries are sampled from the general data for instruction tuning of ChemDFM-R and $\sim$8% are from AM-Deepseek-R1-Distill-1.4M (zhao20251). The ChemDFM-I model is finetuned for one epoch on this mixed dataset.
@@
After distillation, reinforcement learning (RL) is leveraged to further enhance the reasoning capabilities of our model. The composition of the RL dataset is illustrated in Figure 2, while more details of our reinforcement learning process are illustrated in Appendix D and E.